96​% of Dr.​ William's patients end up with​ 20-30 vision or better. Find the probability that exactly 4 of her next 5 patients end up with​ 20-30 vision or better.
Answer #1

Solution :

Given that ,

p = 0.96

1 - p = 1 -0.96 = 0.04  

n = 5

X = 4

Using binomial probability formula ,

P(X = x) = ((n! / (n - x)!) * px * (1 - p)n - x

P(X = 4) =((5! / (5 - 4)!) * 0.964 * 0.045 - 4

= 0.1699

Probability =0.1699